jose is three times as old as alice. Six years hence he will be twice as old as she. How old is each?

Answer by TimothyLamb(4379)

j = 3a
j + 6 = 2(a + 6)
j + 6 = 2a + 12
3a + 6 = 2a + 12
---
a = 6
j = 18
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php